# FlowPolicy: Enabling Fast and Robust 3D Flow-based Policy via Consistency Flow Matching for Robot Manipulation

### v1

@@ Introduction
As illustrated in Figure 3, we selected two typical qualitative examples in Adroit and Metaworld, generated by DP3 and FlowPolicy, respectively. From left to right, this sequence delineates the holistic progression of the task’s execution, with each frame of the sequence corresponding to a distinct phase in the operation, as performed by the agent. On the left is a pen-turning task involving complex dexterity manipulation, which requires turning the blue-capped pen in the dexterous hand to a specific position (i.e., the green-capped pen serves as the reference angle.). Although DP3 accomplishes the dexterity task, the diffusion policy generated based on DP3 fails to ensure consistency with the target pen in 3D space compared to FlowPolicy. Specifically, DP3 turns one end of the pen to a state so high that it is darker in the camera view. In contrast, FlowPolicy generates higher-quality action sequences to turn the pen to the most appropriate angle. More importantly, FlowPolicy gets closer to the target position faster at the same frame. The right case is a difficult ‘Pick-Place’ task in Metaworld, which requires the use of parallel gripping clips to carry a red cube to a blue target location. DP3 unsuccessfully picks up the red cube and fails the task. Instead, FlowPolicy correctly locates the position of the red cube and successfully carries it to the target location. The above findings show that our FlowPolicy not only accomplishes the tasks that DP3 cannot accomplish well, but also achieves a higher quality and speed of accomplishment than DP3. This further demonstrates that our proposed FlowPolicy has a more delicate understanding and generative ability in robot manipulation tasks.
@@
Similarly, we selected four representative tasks in Adroit and Metaworld and visualized their learning curves, as illustrated in Figure 4. Although we defined 3000 epochs for each task, most of the tasks converged and leveled off at 500–1000 epochs. For the simple-level task ‘button press wall’, all three models learned the correct policy easily and converged with a small number of iterations. Although Simple DP3 and DP3 have high success rates, their success rates fluctuate as training progresses. On the contrary, our FlowPolicy learns the button policy well and maintains it, demonstrating the stability of FlowPolicy. For another simple level task ’reach wall’ and a medium difficulty task ‘hammer’, all three models learn faster, but Simple DP3 and DP3 reach a performance bottleneck, while our method achieves a a higher task success rate. The learning curves for the dexterity task in Adroit show similar conclusions. The above results show that FlowPolicy can learn policies faster and better, and demonstrates certain stability.

### v2

@@ Introduction
As illustrated in Figure 3, we selected two typical qualitative examples in Adroit and Metaworld, generated by DP3 and FlowPolicy, respectively. From left to right, this sequence delineates the holistic progression of the task’s execution, with each frame of the sequence corresponding to a distinct phase in the operation, as performed by the agent. On the left is a ‘Pen-turning’ task involving complex dexterity manipulation, which requires turning the blue-capped pen in the dexterous hand to a specific position (i.e., the green-capped pen serves as the reference angle). Although DP3 accomplishes the dexterity task, the diffusion policy generated based on DP3 fails to ensure consistency with the target pen in 3D space compared to FlowPolicy. Specifically, DP3 turns one end of the pen to a state so high that it is darker in the camera view. In contrast, FlowPolicy generates higher-quality action sequences to turn the pen to the most appropriate angle. More importantly, FlowPolicy gets closer to the target position faster at the same frame. The right case is a difficult ‘Pick-Place’ task in Metaworld, which requires the use of parallel gripping clips to carry a red cube to a blue target location. DP3 unsuccessfully picks up the red cube and fails the task. Instead, FlowPolicy correctly locates the position of the red cube and successfully carries it to the target location. The above findings show that our FlowPolicy not only accomplishes the tasks that DP3 cannot accomplish well, but also achieves a higher quality and speed of accomplishment than DP3. This further demonstrates that our proposed FlowPolicy has a more delicate understanding and generative ability in robot manipulation tasks.
@@
Similarly, we selected four representative tasks in Adroit and Metaworld and visualized their learning curves, as illustrated in Figure 4. Although we defined 3000 epochs for each task, most of the tasks converged and leveled off at 500–1000 epochs. For the simple-level task ‘Button Press Wall’, all three models learned the correct policy easily and converged with a small number of iterations. Although Simple DP3 and DP3 have high success rates, their success rates fluctuate as training progresses. On the contrary, our FlowPolicy learns the button policy well and maintains it, demonstrating the stability of FlowPolicy. For another simple level task ’Reach Wall’ and a medium difficulty task ‘Hammer’, all three models learn faster, but Simple DP3 and DP3 reach a performance bottleneck, while our method achieves a higher task success rate. The learning curves for the dexterity task in Adroit show similar conclusions. The above results show that FlowPolicy can learn policies faster and better, and demonstrates certain stability.
@@ Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (NSFC) under Grant Nos. 62372091, 62071097 and in part by Sichuan Science and Technology Program under Grant Nos. 2023NSFSC0462, 2023NSFSC0458, 2023NSFSC1972.